Find values of m so that the function y = x is a solution of the given differential equation.

2y′′ + 9y′ − 7y = 0


1
Expert's answer
2021-11-01T12:35:51-0400
"y=x^m"

"y'=mx^{m-1}"

"y''=m(m-1)x^{m-2}"

Given the differential equation


"x^2y''+9xy'-7y=0"

Substitute


"m(m-1)x^{m-2}x^2+9mx^{m-1}x-7x^m=0"

"m^2-m+9m-7=0"

"m^2+8m-7=0"

"m=-4\pm\sqrt{(4)^2-1(-7)}"

"m_1=-4-\sqrt{23}, m_2=-4+\sqrt{23}"